OpenStudy (anonymous):

elimination method 0.3x-0.2y=4 0.2x+0.5y= -21/23

OpenStudy (anonymous):

gw2011 this one is a little different elimination method 0.3x-0.2y=4 0.2x+0.5y= -21/23

OpenStudy (anonymous):

gw2011 can you help on this one elimination method 0.3x-0.2y=4 0.2x+0.5y= -21/23

OpenStudy (anonymous):

multiply the first equation with 2, multiply the second equation with 3, then subtract the first equation from the second, it will eliminate x, then solve for y , substitute y in any one of the equation to get x